OpenStudy (anonymous):

please help, \[14C _{5} + 7!- 17!/3!14!\] the answer is 6362 but i keep getting a different answer on my calculator

OpenStudy (anonymous):

the 14 should be small like the 5

OpenStudy (asnaseer):

ok - lets look at this one term at a time. what do you get for \(_{14}C_5\)?

OpenStudy (anonymous):

i get 2002

OpenStudy (asnaseer):

that is correct. and 7! ?

OpenStudy (anonymous):

i have a ti-84 and i use the nCr

OpenStudy (anonymous):

5040

OpenStudy (asnaseer):

also correct. so you must have made a mistake in the last term - what did you get for that?

OpenStudy (anonymous):

5.16803703E24

OpenStudy (asnaseer):

that is definitely wrong

OpenStudy (asnaseer):

\[\frac{17!}{3!14!}=\frac{17*16*15*\cancel{14!}}{3!*\cancel{14!}}=\frac{17*16*15}{3*2}\]

OpenStudy (anonymous):

ohh ok i was just putting it straight in my calculator so ill try it like that and let you know if i get it right

OpenStudy (asnaseer):

it looks like what you actually calculated was:\[\frac{17!}{3!}\times14!\]

OpenStudy (asnaseer):

does your calc have brackets? i.e. can you enter '(' and ')'?

OpenStudy (asnaseer):

if it does, then you can try:\[17!\div(3!\times4!)\]

OpenStudy (asnaseer):

the 4! should be 14! up there

OpenStudy (anonymous):

yes my calculator does. i calculated the whole equation separately but still did not get the answer

OpenStudy (asnaseer):

2002 + 5040 - 680 = 6362

OpenStudy (anonymous):

i didn't get the 680 i got 2720

OpenStudy (asnaseer):

use the brackets as I suggested and recalculate it please

OpenStudy (anonymous):

just got it! lol I guess it didn't work before because i did it as 14nCr +7! (17!/3!4!) thank you i got it now!

OpenStudy (asnaseer):

you need to be careful with equations like:\[\frac{ab}{cde}\]when using a calculator. always enter this as:\[a\times b\div (c\times d\times e)\]

OpenStudy (anonymous):

would i be able to do (ab/(cde))?

OpenStudy (asnaseer):

yes - that would also work.

OpenStudy (anonymous):

thank you for your help :)

OpenStudy (asnaseer):

yw :)

OpenStudy (anonymous):

would you happen to know about sequences?

OpenStudy (asnaseer):

I do but unfortunately it is very late here and I need to get some sleep now. Just post it as a new question and I'm sure there will be plenty of other people here willing to help you out.

OpenStudy (anonymous):

ohh ok thank you! good night!

OpenStudy (asnaseer):

thx - good night!